Circular No. 6376

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996V IN NGC 3644 Jean Mueller reports her discovery of a supernova on a IV-N plate

taken on Mar. 28 UT by K. M. Rykoski and Mueller with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survy. SN 1996V (red mag ∼ 16.5 on Apr. 10) is offset 25′′ west and 5′′ north of the center of NGC 3644 (α = 11^{h}21^{m}.5, δ = +2^{o}49′, equinox 2000.0). SN 1996V is not present on a plate taken on 1991 Jan. 15 by C. Brewer and J. D. Mendenhall. An échelle spectrogram (range 400–540 nm, resolution 0.7 nm) taken on Apr. 10 at the Palomar 1.5-m telescope by M. Malkan and H. Teplitz shows broad ripples in the spectrum, including prominent peaks at 471, 502, and 535 nm, consistent with a supernova identification.

COMET C/1995 O1 (HALE-BOPP) H. A. Weaver, Applied Research Corporation; P. D. Feldman, Johns

Hopkins University; M. F. A’Hearn, University of Maryland; C. Arpigny, University of Liège; and J. C. Brandt and C. E. Randall, University of Colorado, write: “During HST spectroscopic observations on Apr. 7 (r = 4.8 AU), we clearly detected OH emission in the A-X (0,0) band near 309 nm with a brightness of ∼ 25 rayleighs in an aperture of 1′′.29 × 3′′.66. Assuming that the observed OH was produced during the photolysis of water that sublimed from the nucleus, we derive a water-production rate of 1.5 × 10^{28} molecules/s. This latter value is only slightly below the CO-production rates derived from submillimeter observations, indicating that water sublimation may soon take over as the primary source of activity in the comet. For a nonrotating spherical nucleus at the comet’s heliocentric distance, we estimate that the maximum sublimation rate cannot exceed the subsolar value of ∼ 9 × 10^{15} molecules cm^{−2} s^{−1}, which implies (assuming negligible emission from the nightside) a lower limit of 10 km for the effective diameter. If the observed OH is actually due to sublimation of water from icy grains in the coma, then the latter conclusion is no longer valid.”

GRO J1744–28 The astrometric investigation by Morrison and Röser on IAUC 6375

was carried out in the absence of their knowledge of IAUC 6369, on which it was shown that the true GRO J1744–28 is located 1′.5 from the candidates that were previously being considered.

1996 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green